### v3

@@ 4.2. Tropical invariant and $F$-invariant
By applying the semifield homomorphism $\beta_{u^{\prime}}$ to $u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})\in\mathbb{F}_{>0}$, we see that
$\beta_{u^{\prime}}(u)=\beta_{u^{\prime}}({\bf x}_{t}^{{\bf g}_{u}^{t}})+\beta_{u^{\prime}}(F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})).$
Since $\beta_{u^{\prime}}({\bf x}_{t})=({\bf a}^{t})^{T}=(\Lambda_{t}{\bf g}_{u^{\prime}}^{t})^{T}$, we have
$\beta_{u^{\prime}}({\bf x}_{t}^{{\bf g}_{u}^{t}})=(\Lambda_{t}{\bf g}_{u^{\prime}}^{t})^{T}{\bf g}_{u}^{t}=[(\Lambda_{t}{\bf g}_{u^{\prime}}^{t})^{T}{\bf g}_{u}^{t}]^{T}=({\bf g}_{u}^{t})^{T}\Lambda_{t}{\bf g}_{u^{\prime}}^{t}.$
We write $F_{u}^{t}=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$, then we know that
$F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{t}^{\widetilde{B}_{t}{\bf v}}.$
We see that
$\beta_{u^{\prime}}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{t}^{\widetilde{B}_{t}{\bf v}})$ $=$ $\max\{({\bf a}^{t})^{T}\widetilde{B}_{t}{\bf v}\mid c_{\bf v}\neq 0\}=\max\{[({\bf a}^{t})^{T}\widetilde{B}_{t}{\bf v}]^{T}\mid c_{\bf v}\neq 0\}$
$=$ $\max\{{\bf v}^{T}\widetilde{B}_{t}^{T}{\bf a}^{t}\mid c_{\bf v}\neq 0\}=\max\{{\bf v}^{T}\widetilde{B}_{t}^{T}\Lambda_{t}{\bf g}_{u^{\prime}}^{t}\mid c_{\bf v}\neq 0\}$
$=$ $\max\{{\bf v}^{T}(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}\mid c_{\bf v}\neq 0\}=F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}].$
Hence, $\beta_{u^{\prime}}(F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}))=\beta_{u^{\prime}}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{t}^{\widetilde{B}_{t}{\bf v}})=F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]$ and
$\beta_{u^{\prime}}(u)=\beta_{u^{\prime}}({\bf x}_{t}^{{\bf g}_{u}^{t}})+\beta_{u^{\prime}}(F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}))=({\bf g}_{u}^{t})^{T}\Lambda_{t}{\bf g}_{u^{\prime}}^{t}+F_{u}^{t}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t}]=\langle u,u^{\prime}\rangle_{t}.$
Since the semifield homomorphism $\beta_{u^{\prime}}$ is uniquely determined by $u^{\prime}$, we know that the integer $\beta_{u^{\prime}}(u)$ only depends on $u$ and $u^{\prime}$, not on the choice of an initial seed. The equality $\beta_{u^{\prime}}(u)=\langle u,u^{\prime}\rangle_{t}$ says that each choice of a vertex $t\in\mathbb{T}_{n}$ provides an explicit formula to calculate the integer $\beta_{u^{\prime}}(u)$. This completes the proof. ∎
Remark 4.9.
From the proof above, we see that the integer $\langle u,u^{\prime}\rangle_{w}$ can be interpreted as the evaluation of the positive function $u\in\mathbb{F}_{>0}$ (on cluster $\mathscr{A}$-variety in the sense of [FG-2009]) at the tropical point $[{\bf a}]:=\{{\bf a}^{t}=\Lambda_{t}{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ (associated to the same cluster $\mathscr{A}$-variety) determined by $u^{\prime}$.
Since the tropical invariant $\langle u,u^{\prime}\rangle_{w}$ only depends on $u$ and $u^{\prime}$, not on the choice of $w\in\mathbb{T}_{n}$, we denote by $\langle u,u^{\prime}\rangle:=\langle u,u^{\prime}\rangle_{w}$ and call $\langle-,-\rangle:\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}$ the tropical pairing of $\mathcal{U}$.
Definition 4.10 ($F$-invariant and $F$-compatibility).
Let $u$ and $u^{\prime}$ be two good elements in $\mathcal{U}$.
  • (i)
    The $F$-invariant $(u\mid\mid u^{\prime})_{F}$ between $u$ and $u^{\prime}$ is defined to be the symmetrized sum of the tropical invariants:
    $(u\mid\mid u^{\prime})_{F}=\langle u,u^{\prime}\rangle+\langle u^{\prime},u\rangle=\langle u,u^{\prime}\rangle_{w}+\langle u^{\prime},u\rangle_{w},$
    where $w$ is any vertex in $\mathbb{T}_{n}$.
  • (ii)
    $u$ and $u^{\prime}$ are said to be $F$-compatible, if $(u\mid\mid u^{\prime})_{F}=0$.
Proposition 4.11.
Let $u$ and $u^{\prime}$ be two good elements in $\mathcal{U}$. The following statements hold.
  • (i)
    $(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]$, where $w$ is any vertex in $\mathbb{T}_{n}$.
  • (ii)
    $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
  • (iii)
    If $u$ is a cluster monomial, then $(u\mid\mid u)_{F}=0$.
Proof.
(i) Let $w$ be a vertex in $\mathbb{T}_{n}$. By the definition of $(u\mid\mid u^{\prime})_{F}$, we see that
$(u\mid\mid u^{\prime})_{F}$ $=$ $\langle u,u^{\prime}\rangle_{w}+\langle u^{\prime},u\rangle_{w}$
@@
(ii) Let $w$ be a vertex in $\mathbb{T}_{n}$. Since both $F_{u}^{w}$ and $F_{u^{\prime}}^{w}$ are polynomials with constant term $1$, we have
$F_{u}^{w}[{\bf h}],\;F_{u^{\prime}}^{w}[{\bf h}]\in\mathbb{Z}_{\geq 0}$
for any ${\bf h}\in\mathbb{Z}^{n}$. Then by (i), we have $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
(iii) Say $u$ is a cluster monomial in seed $({{\bf x}_{w},\widetilde{B}_{w}})$. Then $F_{u}^{w}=1$, where $F_{u}^{w}$ is the $F$-polynomial of $u$ with respect to vertex $w\in\mathbb{T}_{n}$. Then by (i), we see that $(u\mid\mid u)_{F}=0$. ∎
@@
(i) The results in Proposition 4.11 (iii) are analogous to [DWZ10, Corollary 7.2] for the $E$-invariant in the additive categorification of cluster algebras and to [kkko-2018, Lemma 3.2.3] for the $\mathfrak{d}$-invariant in the monoidal categorification of cluster algebras.
(ii) The good elements satisfying $(u\mid\mid u)_{F}=0$ are analogous to the rigid objects in cluster categories and to the real simples in the monoidal categorification of cluster algebras.
@@
Recently, Fei introduced and studied the tropical $F$-polynomials in [fei_2019a, fei_2019b] from the viewpoint of representation theory. One can see that the $F$-invariant is the sum of the evaluations of two tropical $F$-polynomials. When restricting to cluster monomials, [fei_2019b, Theorem 3.22] implies that the $F$-invariant is essentially the same as the $E$-invariant in $\tau$-tilting theory. I will give more details about this in [cao_2025].
Remark 4.14.
(i) If we take $u=x_{i;t}$ and $u^{\prime}=x_{j;t}$, then ${\bf g}_{u}^{t}={\bf e}_{i},\;{\bf g}_{u^{\prime}}^{t}={\bf e}_{j}$ and $F_{u}^{t}=1$, where ${\bf e}_{k}$ is the $k$th column of $I_{m}$ for $k\in[1,m]$. By the definition of tropical invariant, we have
$\langle u,u^{\prime}\rangle=\langle u,u^{\prime}\rangle_{t}=\langle x_{i;t},x_{j;t}\rangle_{t}={\bf e}_{i}^{T}\Lambda_{t}{\bf e}_{j}+0=\lambda_{ij;t},$
where $\lambda_{ij;t}$ is the $(i,j)$-entry of $\Lambda_{t}$. So the tropical invariant $\langle x_{i;t},x_{j;t}\rangle$ is related to the quantization matrix $\Lambda_{t}$. This is analogous to [kkko-2018, Proposition 7.1.2 (a)] for the $\Lambda$-invariant $\Lambda(M,N)$ in the monoidal categorification of cluster algebras.
(ii) Notice that the $F$-invariant $(u\mid\mid u^{\prime})_{F}=\langle u,u^{\prime}\rangle+\langle u^{\prime},u\rangle$ in cluster algebras is defined to be the symmetrized sum of the tropical invariants while the $\mathfrak{d}$-invariant $\mathfrak{d}(M,N)=\frac{1}{2}(\Lambda(M,N)+\Lambda(N,M))$ in the monoidal categorification [kkko-2018] is defined to be half of the symmetrized sum of $\Lambda$-invariants. This suggests that the $F$-invariant in cluster algebras is related with the $\mathfrak{d}$-invariant in the monoidal categorification.
@@
We know that $F_{x_{i;t}}^{t}=1$ and ${\bf g}_{x_{i;t}}^{t}={\bf e}_{i}$, where ${\bf e}_{i}$ is the $i$th column of $I_{m}$. Now we use the vertex $t$ to calculate $(x_{i;t}\mid\mid u)_{F}$ and by using the formula in Proposition 4.11 (i), we have
@@
As we said that $F$-invariant can be viewed as an analog of the $E$-invariant in the additive categorification of cluster algebras and the $\mathfrak{d}$-invariant in the monoidal categorification of cluster algebras. We provide more reasons in the following remark.
@@
(i) By Proposition 4.15, we know that the $F$-invariant $(x_{i;t}\mid\mid u)_{F}$ is related to the components of $f$-vectors.
(ii) By the equality (1.1) in the introduction part, the $E$-invariant is related to the extension dimension in the cluster categories. Such dimensions are used to categorify the components of $f$-vectors.
(iii) The components of $f$-vectors measure the distance between the highest degree of some $y_{k}$ and the lowest degree of $y_{k}$ in $F$-polynomials, while the $\mathfrak{d}$-invariant has the similar meaning, cf., [kkko-2018, Corollary 4.2.4] and [qin-2020, Lemma 5.1].
Remark 4.17.
(i) The $F$-invariant $(-\mid\mid-)_{F}:\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}_{\geq 0}$ is a great improvement of the $f$-compatibility degree introduced by Fu and Gyoda in [Fu-Gyoda]. The $f$-compatibility degree, defined using the components of $f$-vectors of cluster variables, is a function
$(-\mid\mid-)_{f}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq 0},$
where $\mathcal{X}$ is the set of unfrozen cluster variables in $\mathcal{U}$. When restriced to $\mathcal{X}$, the $F$-invariant and the $f$-compatibility degree are related as follows:
@@
    Let $k$ be an integer in $[1,n]$ and let $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$. Suppose that $(x_{k;t}\mid\mid u)_{F}\neq 0$ but $(x_{i;t}\mid\mid u)_{F}=0$ for any $i\in[1,n]\backslash\{k\}$. Then $(x_{i;t^{\prime}}\mid\mid u)_{F}=0$ for any $i\in[1,n]$. In particular, $u$ is a cluster monomial in ${\bf x}_{t^{\prime}}$.
@@
Let $u$ be a $[{\bf g}]$-bigood element in $\mathcal{U}$, and let ${\bf x}_{t}^{\bf h}=\prod_{j=1}^{m}x_{j;t}^{h_{j}}$ be a cluster monomial of $\mathcal{U}$, i.e., ${\bf h}^{\circ}:=(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$. Then
@@
where ${\bf g}_{u}^{t}$ is the value of $[{\bf g}]$ at vertex $t$. Since $u$ is a $[{\bf g}]$-bigood element, $F_{u}^{t}$ has the form:
@@
where ${\bf f}_{u}^{t}=(f_{1}^{t},\ldots,f_{n}^{t})^{T}$ is the $f$-vector of $u$ with respect to vertex $t$. Since ${\bf x}_{t}^{\bf h}$ is a cluster momomial, we know that $F_{{\bf x}_{t}^{\bf h}}^{t}=1$ and $(h_{1},\ldots,h_{n})\in\mathbb{Z}_{\geq 0}^{n}$. By using the chart ${\bf x}_{t}$ to calculate $({\bf x}_{t}^{\bf h}\mid\mid u)_{F}$, we have
@@
(The last equality comes from the facts that $F_{{\bf x}_{t}^{\bf h}}^{w}=\prod_{j=1}^{m}(F_{x_{j;t}}^{w})^{h_{j}}=\prod_{j=1}^{n}(F_{x_{j;t}}^{w})^{h_{j}}$ and that for any two polynomials $F_{1},F_{2}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ and any vector ${\bf r}\in\mathbb{Z}^{n}$, we have
@@
by the definition of tropical polynomials.) On the other hand, we have
@@
$[{\bf g}_{z}]:=\{{\bf g}_{z}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$
corresponding to the cluster variable $z$. Since $(x\mid\mid z)_{F}=0$ and by Corollary 4.18 (i), we know that the $k$th component $g_{k;z}^{t}$ of ${\bf g}_{z}^{t}$ satisfies $g_{k;z}^{t}\geq 0$ whenever $x=x_{k;t}$.
On the other hand, by [cao-li-2020, Theorem 8], we know that there exists a seed $({\bf x}_{w},\widetilde{B}_{w})$ containing $x$ (say $x=x_{k;w}$) such that $g_{i;z}^{w}\geq 0$ for any $i\in[1,n]$ with $x_{i;t}\neq x$, i.e., for any $i\in[1,n]\setminus\{k\}$. Thus for the vertex $w$, we have $g_{j;z}^{w}\geq 0$ for $j=1,\ldots,n$, i.e., the principal part $({\bf g}_{z}^{t})^{\circ}$ of ${\bf g}_{z}^{t}$ is in $\mathbb{Z}_{\geq 0}^{n}$. Thus $u:={\bf x}_{w}^{{\bf g}_{z}^{w}}$ is a cluster monomial of $\mathcal{U}$. Hence, the two cluster monomials $u={\bf x}_{w}^{{\bf g}_{z}^{w}}$ and $z$ have the same extended $g$-vector ${\bf g}_{z}^{w}$ with respect to vertex $w$. It is well-known that different cluster monomials have different extended $g$-vectors, c.f., [GHKK18]. So we must have $z=u={\bf x}_{w}^{{\bf g}_{z}^{w}}$. This implies $z$ is a cluster variable in ${\bf x}_{w}$. In particular, $z$ and $x=x_{k;w}$ are contained in the same cluster ${\bf x}_{w}$. ∎
@@
Let continue with Example 3.8. Take $\Lambda=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}=\widetilde{B}$. We know that $\widetilde{B}^{T}\Lambda=(S\mid{\bf 0})=I_{2}$ and the canonical expressions of $x_{3},x_{4},x_{5}$ with respect to the initial seed $t_{0}=({\bf x},\widetilde{B})$ are given as follows:
@@
Since $u$ and $u^{\prime}$ are cluster monomials and by Proposition 4.11 (iii), we have
@@
Two tropical points $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and $[{\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ are said to be sign-coherent, if for any vertex $t\in\mathbb{T}_{n}$ and any $k\in[1,n]$, we have $g_{k}^{t}\cdot(g_{k}^{\prime})^{t}\geq 0$, where $g_{k}^{t}$ and $(g_{k}^{\prime})^{t}$ are the $k$th components of ${\bf g}^{t}$ and $({\bf g}^{\prime})^{t}$.
@@
Let $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and $[{\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ be two tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. If $[{\bf g}]$ and $[{\bf g}^{\prime}]$ are sign-coherent, then the collection $\{{\bf g}^{t}+({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is still a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. We denote it by $[{\bf g}\uplus{\bf g}^{\prime}]$.
@@
Let $u$ be a $[{\bf g}]$-good element in $\mathcal{U}$ and $u^{\prime}$ a $[{\bf g}^{\prime}]$-good element in $\mathcal{U}$. Suppose that $(u\mid\mid u^{\prime})_{F}=0$. Then the following statements hold.
@@
    The two tropical points $[{\bf g}]$ and $[{\bf g}^{\prime}]$ are sign-coherent. In particular, the tropical point $[{\bf g}\uplus{\bf g}^{\prime}]$ can be defined.
@@
    The product $uu^{\prime}$ is a $[{\bf g}\uplus{\bf g}^{\prime}]$-good element in $\mathcal{U}$.

### v4

@@ 4.2. Tropical invariant and $F$-invariant
such that $\beta_{u^{\prime}}({\bf x}_{t})=(\Lambda_{t}{\bf g}_{u^{\prime}}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. By applying the semifield homomorphism $\beta_{u^{\prime}}$ to $u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\widehat{\bf y}_{w})\in\mathbb{F}_{>0}$, we have
$\beta_{u^{\prime}}(u)=\beta_{u^{\prime}}({\bf x}_{w}^{{\bf g}_{u}^{w}})+\beta_{u^{\prime}}(F_{u}^{w}(\widehat{\bf y}_{w})).$
Since $\beta_{u^{\prime}}({\bf x}_{w})=(\Lambda_{w}{\bf g}_{u^{\prime}}^{w})^{T}$, we have
$\beta_{u^{\prime}}({\bf x}_{w}^{\bf h})=(\Lambda_{w}{\bf g}_{u^{\prime}}^{w})^{T}{\bf h}=[(\Lambda_{w}{\bf g}_{u^{\prime}}^{w})^{T}{\bf h}]^{T}={\bf h}^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w},\quad\forall{\bf h}\in\mathbb{Z}^{m}.$
We write $F_{u}^{w}({\bf y})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}\in\mathbb{N}[y_{1},\ldots,y_{n}]$. Then $F_{u}^{w}(\widehat{\bf y}_{w})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{w}^{\widetilde{B}_{w}{\bf v}}$. Thus
$\beta_{u^{\prime}}(F_{u}^{w}(\widehat{\bf y}_{w}))$ $=\beta_{u^{\prime}}(\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf x}_{w}^{\widetilde{B}_{w}{\bf v}})=\max\{{\bf v}^{T}\widetilde{B}_{w}^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}\mid c_{\bf v}\neq 0\}$
$=\max\{{\bf v}^{T}(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}\mid c_{\bf v}\neq 0\}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}].$
Hence, we have
$\langle u,u^{\prime}\rangle_{\rm trop}=\beta_{u^{\prime}}(u)=\beta_{u^{\prime}}({\bf x}_{w}^{{\bf g}_{u}^{w}})+\beta_{u^{\prime}}(F_{u}^{w}(\widehat{\bf y}_{w}))=({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}].$
(ii) By (i) and the definition of $F$-invariant, we have
$(u\mid\mid u^{\prime})_{F}$ $=\langle u,u^{\prime}\rangle_{\rm trop}+\langle u^{\prime},u\rangle_{\rm trop}$
@@
Since both $F_{u}^{w}$ and $F_{u^{\prime}}^{w}$ are polynomials with constant term $1$, we have
$F_{u}^{w}[{\bf r}],\;F_{u^{\prime}}^{w}[{\bf r}]\;\in\mathbb{Z}_{\geq 0}$
for any ${\bf r}\in\mathbb{Z}^{n}$. So $(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]\;\in\mathbb{Z}_{\geq 0}$.
(iii) Say $u$ is a cluster monomial in seed $({{\bf x}_{w},\widetilde{B}_{w}})$. Then $F_{u}^{w}=1$. Then by (ii), we see that $(u\mid\mid u)_{F}=0$. ∎
Since the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ can be written as
$(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}],$
we call the non-negative integer $F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]$ the partial $F$-invariant of $(u,u^{\prime})$ at vertex $w\in\mathbb{T}_{n}$.
@@
Notice that $(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}=S({\bf g}_{u^{\prime}}^{w})^{\circ}$, where $({\bf g}_{u^{\prime}}^{w})^{\circ}\in\mathbb{Z}^{n}$ is the principal part of ${\bf g}_{u}^{w}\in\mathbb{Z}^{m}$. This suggests that the notion of partial $F$-invariant at vertex $w\in\mathbb{T}_{n}$ and the $F$-invariant in some sense only depend on the principal part of the mutation matrices (i.e., the exchange matrices) of $\mathcal{A}$ and a choice of the skew-symmetrizer $S$.
@@
Recently, Fei introduced and studied the tropical $F$-polynomials in [fei_2019a, fei_2019b] from the viewpoint of representation theory of finite-dimensional algebras and [fei_2019b, Theorem 3.22] suggests that the partial $F$-invariant coincides with the partial $E$-invariant for Jacobi-finite quivers with potentials.
@@
We know that $F_{x_{i;t}}^{t}=1$ and ${\bf g}_{x_{i;t}}^{t}={\bf e}_{i}$, where ${\bf e}_{i}$ is the $i$th column of $I_{m}$. Now we use the vertex $t$ to calculate $(x_{i;t}\mid\mid u)_{F}$. By using the formula in Theorem 4.10 (ii), we have
@@
(i) The $F$-invariant $(-\mid\mid-)_{F}\colon\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}_{\geq 0}$ is a great improvement of the $f$-compatibility degree introduced by Fu and Gyoda in [Fu-Gyoda] and thus the compatibility degree introduced by Fomin and Zelevinsky in [fz-2003y]. The $f$-compatibility degree, defined using the components of $f$-vectors of cluster variables, is a function
$(-\mid\mid-)_{f}\colon\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq 0},$
where $\mathcal{X}$ is the set of unfrozen cluster variables in $\mathcal{U}$. By Proposition 4.13, when restricted to $\mathcal{X}$, the $F$-invariant and the $f$-compatibility degree are related as follows:
@@
[…]
@@
Let $u$ be a $\bf g$}]-bigood element in $\mathcal{U}$, and let ${\bf x}_{t}^{\bf h}=\prod_{j=1}^{m}x_{j;t}^{h_{j}}$ be a cluster monomial of $\mathcal{U}$, i.e., ${\bf h}^{\circ}\coloneqq(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$. Then
@@
where ${\bf g}_{u}^{t}$ is the value of $\bf g$}] at vertex $t$. Since $u$ is a $\bf g$}]-bigood element, $F_{u}^{t}$ has the form:
@@
where ${\bf f}_{u}^{t}=(f_{1}^{t},\ldots,f_{n}^{t})^{T}$ is the $f$-vector of $u$ with respect to vertex $t$. Since ${\bf x}_{t}^{\bf h}$ is a cluster momomial, we know that $F_{{\bf x}_{t}^{\bf h}}^{t}=1$ and $(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$. By using the cluster ${\bf x}_{t}$ to calculate $({\bf x}_{t}^{\bf h}\mid\mid u)_{F}$, we have
@@
The last equality comes from the facts that $F_{{\bf x}_{t}^{\bf h}}^{w}=\prod_{j=1}^{m}(F_{x_{j;t}}^{w})^{h_{j}}=\prod_{j=1}^{n}(F_{x_{j;t}}^{w})^{h_{j}}$ and that for any two polynomials $F_{1},F_{2}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ and any vector ${\bf r}\in\mathbb{Z}^{n}$, we have
@@
by Proposition 4.9 (ii).
On the other hand, we know that
@@
${\bf g}_{z}]\coloneqq\{{\bf g}_{z}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$
corresponding to the cluster variable $z$. Since $(x\mid\mid z)_{F}=0$ and by Corollary 4.15 (i), we know that the $k$th component $g_{k;z}^{t}$ of ${\bf g}_{z}^{t}$ satisfies $g_{k;z}^{t}\geq 0$ whenever $x=x_{k;t}$ for a vertex $t$.
On the other hand, by [cao-li-2020, Theorem 8], we know that there exists a seed $({\bf x}_{w},\widetilde{B}_{w})$ containing $x$ (say $x=x_{k;w}$) such that $g_{i;z}^{w}\geq 0$ for any $i\in[1,n]$ with $x_{i;t}\neq x$, i.e., for any $i\in[1,n]\setminus\{k\}$. Thus for the vertex $w$, we have $g_{j;z}^{w}\geq 0$ for $j=1,\ldots,n$, i.e., the principal part $({\bf g}_{z}^{t})^{\circ}$ of ${\bf g}_{z}^{t}$ belongs to $\mathbb{Z}_{\geq 0}^{n}$. Thus ${\bf x}_{w}^{{\bf g}_{z}^{w}}$ is a cluster monomial of $\mathcal{U}$. Hence, the two cluster monomials ${\bf x}_{w}^{{\bf g}_{z}^{w}}$ and $z$ have the same extended $g$-vector ${\bf g}_{z}^{w}$ with respect to vertex $w$. It is well-known that different cluster monomials have different extended $g$-vectors, cf. [GHKK18]. So we must have $z={\bf x}_{w}^{{\bf g}_{z}^{w}}$. This implies that $z$ is a cluster variable in ${\bf x}_{w}$. In particular, $z$ and $x=x_{k;w}$ are contained in the same cluster ${\bf x}_{w}$. ∎
@@
[…]
@@
Since $u$ and $u^{\prime}$ are cluster monomials and by Theorem 4.10 (iii), we have
@@
Two tropical points ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and ${\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ are said to be sign-coherent, if for any vertex $t\in\mathbb{T}_{n}$ and any $k\in[1,n]$, we have $g_{k}^{t}\cdot(g_{k}^{\prime})^{t}\geq 0$, where $g_{k}^{t}$ and $(g_{k}^{\prime})^{t}$ are the $k$th components of ${\bf g}^{t}$ and $({\bf g}^{\prime})^{t}$.
@@
Let ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and ${\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ be two tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. If $\bf g$}] and $\bf g$$^{\prime}$] are sign-coherent, then the collection $\{{\bf g}^{t}+({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is still a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. We denote it by ${\bf g}\uplus{\bf g}^{\prime}]$.
@@
Let $u$ be a $\bf g$}]-good element in $\mathcal{U}$ and $u^{\prime}$ a $\bf g$$^{\prime}$]-good element in $\mathcal{U}$. Suppose that $(u\mid\mid u^{\prime})_{F}=0$. Then the following statements hold.
@@
    The two tropical points $\bf g$}] and $\bf g$$^{\prime}$] are sign-coherent. In particular, the tropical point ${\bf g}\uplus{\bf g}^{\prime}]$ can be defined.
@@
    The product $uu^{\prime}$ is a ${\bf g}\uplus{\bf g}^{\prime}]-$good element in $\mathcal{U}$.